OpenStudy (anonymous):

plz check my answer What is the simplified form of 3/x^2/1/x^3 i got 3/x for my answer

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

\[\frac{\frac{3}{x^2}}{\frac{1}{x^3}} = \frac{3}{x^2} \times \frac{x^3}{1} = \frac{3\times x^3}{x^2} = 3x\]

OpenStudy (anonymous):

oh ok thanks i see where i made my mistake.

OpenStudy (austinl):

No problem!